OpenStudy (anonymous):

find the difference quotient of f(x)= (2x)/(x-1)

OpenStudy (anonymous):

@Nnesha can you please help me

Nnesha (nnesha):

we need to use formula f(x+h)-f(x) over h right ?

OpenStudy (anonymous):

instead of x im using a but yes

OpenStudy (anonymous):

f(a)= (2a)/(a-1)

Nnesha (nnesha):

ohh okay so first we have to find f(a+h) substitute a for a+h

OpenStudy (anonymous):

and f(a+h)= (2a+2h)/(-a-h)

OpenStudy (anonymous):

am i correct

Nnesha (nnesha):

the denominator would be a-1 = (a+h)-1 so it would be a+h-1 it's not a multiplication

OpenStudy (anonymous):

oh alright

Nnesha (nnesha):

alright so now we can substitute f(a+h) for that into the formula \[\huge \rm \frac{ \frac{ 2a+2h }{ a+h-1 } +\frac{ 2a }{ a-1 } }{ h }\] now find the common denominator

OpenStudy (anonymous):

so you multiply it by h?

Nnesha (nnesha):

first deal with the numerator 2a+2h/a+h-1 + 2a/a-1

OpenStudy (anonymous):

what about it

Nnesha (nnesha):

what would be the common denominator ?

Nnesha (nnesha):

brb i'll login to other laptop

OpenStudy (anonymous):

shouldnt it be a+h-1?

Nnesha (nnesha):

both denominator are not the same so we should multiply both denominator \[\huge\rm \frac{ (\color{Red}{a-1})(2a+2h) + 2a(\color{blue}{a+h-1}) }{ (a+h-1)(a-1)}\] and when we find common denominator we should multiply numerator of first fraction by the denominator of 2nd fraction

Nnesha (nnesha):

and multiply the numerator of 2nd fraction with the denominator of first

OpenStudy (anonymous):

so what did get

Nnesha (nnesha):

 kinda like cross multiplication